OpenStudy (anonymous):

Find two supplementary angles if the measure of the first angle is three times the measure of the second

OpenStudy (campbell_st):

so let the smaller angle be x the larger angle is 3x then 3x + x = 180 the rule for supplementary angles now solve for x

OpenStudy (anonymous):

Angles are supplementary if when you add up all the angle degrees, they sum up to 180. So, then this problem becomes algebra. One angle, let's call it x, is 3 times as large as the other angle, which we can call y. Since they are supplementary, \[x + y = 180\]. Since x is 3 times as large as y, \[x = 3y\] We can substitute this back in to \[x + y = 180\] to get \[y + 3y = 4y = 180\] dividing both sides by 4, we get \[y = 4\] Since we know the relationship\[x = 3y\] we can substitute y =45 back in to get\[x = 3*(45) = 135\] so x = 135 and y = 45

OpenStudy (anonymous):

Omg thank you so much guys it worked out for me! Is it ok If I message you guys if I get lost again?